OpenStudy (anonymous):

For any three consecutive integers, show that the average of the smallest and largest integers is the middle integer

OpenStudy (mertsj):

x=first integer x+1=next consecutive integer x+2= third consecutive integer. Add them up and divide by 3 You should get x+1 which is the middle integer

OpenStudy (ranga):

Let the 3 consecutive integer be: N-1, N and N+1 The average of the smallest and the largest integer is: ( N -1 + N + 1 ) / 2 = 2N/2 = N which is the middle integer.

OpenStudy (anonymous):

Got it thank you!